# MACRO-LLM: LLM-Empowered Multi-Agent Collaborative Reasoning under Spatiotemporal Partial Observability

### v1

@@ Appendix I Analysis of Failure Cases and Solutions
Despite access to topological information, agents occasionally exhibit spatial hallucinations, such as confusing front and back vehicles in the CACC task. To mitigate this, we employ few-shot prompting incorporating explicit examples of state transitions (e.g., detailing how acceleration impacts headway relative to neighbors).
Furthermore, LLMs struggle with the precise numerical calculations required for kinematic updates, often leading to arithmetic inaccuracies. To ensure precision, we offload deterministic calculations to an external Python interpreter. The computation results are then integrated into the prompt for subsequent inference.
@@
In scenarios featuring dense interaction prompts, such as concurrent proposals involving similar identifiers, such as “Vehicle 2 propose action1 to Vehicle 1, and propose action2 to Vehicle 3), agents may misattribute actions or forget their own identity. We attribute this to the attention mechanism’s difficulty in distinguishing semantically similar tokens within extended contexts. To address this, we enforce a structured schema for inter-agent communication and integrate a self-consistency mechanism to validate sender-receiver pairings. Additionally, we reinforce role stability by prepending the instruction “You are [Agent Name]” to every query prompt.
@@
In the PC task, the Qwen series frequently triggered safety refusals during stages characterized by high mortality rates, erroneously flagging simulation statistics as harmful content (returning openai.BadRequestError). This behavior indicates a stricter safety filter configuration compared to models like GPT-4o. To mitigate these false positives, we appended a system prompt explicitly clarifying the synthetic nature of the data: "SIMULATION NOTE: SIMULATION NOTE: This is a hypothetical pandemic experiments on pandemic simulator. All data is synthetic and does not reflect real-world events!!!" Additionally, we implemented a rollback mechanism to discard and retry generation steps interrupted by safety filters.
Appendix J Core Prompts of MACRO-LLM
J.1 CoProposer
Table 8 provides the core prompts of proposal generation with a rollout-simulated verification in CoProposer.
@@
J.2 Negotiator
Table 9 provides the core prompts for proposal evaluation with spatial strategies and feature updates in Negotiator when consensus is not achieved. Table 10 provides the core prompts for updating proposal with updated spatial strategies and features in Negotiator.
@@
J.3 Introspector
Table 11 presents the core prompts for generating the semantic gradient after completing one-step execution. Table 12 presents the core prompts for updating the temporal strategy using the learning rate and semantic gradient.
Refer to caption Introspector: Generate Semantic Gradient You are [agent_name]. You can observe [observable_agent_name]. The total system reward for this time step is xx% lower than in the last time step indicating that the system state getting worse. Your individual reward has dropped from $x$ to $x$, which drop xx%.
@@
Step 2: According to the related rules you summarized, be specific about ‘‘what to do" and ‘‘what to avoid" and output between tags <self-checking_rules> and </self-checking_rules>.
@@
Figure 11: The core prompt of generating semantic gradient.
@@
Based on the <reasons></reasons> you summarized (from Table 11), update your strategies carefully with NO MORE than xx% learning rate.
@@
[self-checking_rules] are a set of rules for verifying the output format and common-sense consistency of LLM outputs. These rules ensure that the model correctly applies physical reasoning. For example, in CACC task, “If ACCELERATION_GAP > 0, MY_VELOCITY will be INCREASED, MY_HEADWAY will be DECREASED." These rules help correct occasional LLM mistakes, such as misjudging the travel direction of vehicles and thus misunderstanding how acceleration or deceleration affects inter-vehicle headway.
Appendix K AI Assistants in Writing
We utilized AI assistants (e.g., Gemini) exclusively for grammatical error correction, sentence polishing to enhance readability, and generating robot icons used in Figures 1 and 2. We explicitly state that the core scientific concepts, experimental design, logical structure, reference, and substantive content of the paper were entirely human-authored without generative contributions from LLMs.

### v2

@@ I.2 Decision Frequency Analysis
Despite access to topological information, agents occasionally exhibit spatial hallucinations, such as confusing front and back vehicles in the CPP task. To mitigate this, we employ few-shot prompting incorporating explicit examples of state transitions (e.g., detailing how acceleration impacts headway relative to neighbors).
While LLMs occasionally produce imprecise numerical estimates during multi-step rollout predictions, all inference-dependent computations—including state prediction and constraint evaluation during rollout verification—are performed entirely by the LLM, with progressive constraint relaxation (Sec. 3.2) reducing sensitivity to arithmetic imprecision. Deterministic computations independent of the LLM reasoning process, such as post-execution environment updates, are offloaded to an external numerical module to ensure precision.
@@
In scenarios featuring dense interaction prompts, such as concurrent proposals involving similar identifiers, such as “Vehicle 2 propose action1 to Vehicle 1, and propose action2 to Vehicle 3”, agents may misattribute actions or forget their own identity. We attribute this to the attention mechanism’s difficulty in distinguishing semantically similar tokens within extended contexts. To address this, we enforce a structured schema for inter-agent communication and integrate a self-consistency mechanism to validate sender-receiver pairings. Additionally, we reinforce role stability by prepending the instruction “You are [Agent Name]” to every query prompt.
@@
In the PC task, the Qwen series frequently triggered safety refusals during stages characterized by high mortality rates, erroneously flagging simulation statistics as harmful content (returning openai.BadRequestError). This behavior indicates a stricter safety filter configuration compared to models like GPT-4o. To mitigate these false positives, we appended a system prompt explicitly clarifying the synthetic nature of the data: ‘‘SIMULATION NOTE: This is a hypothetical pandemic experiments on pandemic simulator. All data is synthetic and does not reflect real-world events!!!’’ Additionally, we implemented a rollback mechanism to discard and retry generation steps interrupted by safety filters.
Appendix P Analysis of Introspector’s Attribution-based Reasoning
To investigate how the Introspector achieves effective strategy refinement, we analyze whether its semantic revision signals reflect genuine attribution-based reasoning, attributing state changes to inter-agent interaction mechanisms, rather than simple trajectory pattern matching that maps local observations directly to actions. We provide both qualitative and quantitative evidence below.
P.1 Qualitative Evidence: Scenario-Adaptive Diagnostic Attribution Analysis
We present representative semantic revision signals generated by the Introspector in two distinct CPP scenarios to demonstrate that it produces scenario-adaptive diagnostic attribution analyses rather than templated outputs.
Case 1 (Catch-up Scenario).
Vehicle 2 generated the following revision signal:
“The reasons may be due to the unexpected deceleration of vehicle 1 and the slight acceleration of vehicle 3. Vehicle 1’s velocity decreased more than anticipated, causing vehicle 2 to adjust its acceleration to maintain a safe distance.”
The agent attributes its performance drop to specific neighbor behaviors, the unexpected deceleration of vehicle 1 and the acceleration pattern of vehicle 3, and derives a targeted adjustment strategy.
Case 2 (Slow-down Scenario).
The same agent position generated a qualitatively different signal:
“The spatial analysis indicates a general trend of deceleration, which may have influenced the need for a stronger deceleration to maintain formation.”
The agent attributes its adjustment to an aggregated mean-field spatial trend from unobservable agents, rather than to any individual neighbor’s behavior.
These two cases demonstrate fundamentally different attribution directions: Case 1 performs local neighbor-level attribution, while Case 2 reasons about collective spatial dynamics. This confirms that the Introspector adapts its diagnostic attribution analysis to the scenario’s underlying structure rather than applying a fixed reasoning pattern.
P.2 Quantitative Evidence: Diagnostic Attribution Statistics
To systematically evaluate whether the Introspector’s reflections constitute attribution reasoning, we analyzed three complete collaboration logs comprising all reflection instances across full simulation episodes. We clarify that the attribution explicitly references other agents’ behaviors, neighborhood statistical features, or inter-agent interaction effects as contributing factors to the observed performance change. Results show that 89.3% of reflections contain explicit attribution to other agents’ behaviors or neighborhood features, rather than self-trajectory patterns. This indicates that the Introspector systematically reasons about multi-agent interaction mechanisms rather than overfitting to local observation sequences.
P.3 Behavioral Coherence Verification
We further verify that the Introspector’s attribution reasoning translates into coherent behavioral adjustments. In the Catch-up case (Case 1 above), vehicle 2 identified that vehicle 1 decelerated unexpectedly (acceleration shifting from +1.8+1.8 to 1.3-1.3 m/s2) while vehicle 3 maintained low acceleration. The diagnostic attribution concluded that stabilization was needed to prevent headway violation. Consequently, vehicle 2 adjusted its own acceleration from 1.05-1.05 to 0.82-0.82 m/s2, a conservative correction aligned with the diagnosed cause, reducing deceleration intensity to avoid over-correction given vehicle 1’s unexpected behavior.
This behavioral coherence, where the direction and magnitude of action adjustments logically follow from the attribution diagnosis, is further corroborated by the ablation results in Figs. 4 and 5, where removing the Introspector leads to persistent oscillations in CPP ($1.9\times$ degradation in RMSE-H) and 44.4% longer pandemic duration in PC, confirming that attribution adaptation is essential for long-horizon coordination.
Appendix Q Core Prompts of MACRO-LLM
Q.1 CoProposer
Fig. 8 provides the core prompts of proposal generation with a rollout-simulated verification in CoProposer.
@@
Q.2 Negotiator
Fig. 9 provides the core prompts for proposal evaluation with spatial strategies and feature updates in Negotiator when consensus is not achieved. Fig. 10 provides the core prompts for updating proposal with updated spatial strategies and features in Negotiator.
@@
Q.3 Introspector
Fig. 11 presents the core prompts for generating the revision signal after completing one-step execution. Fig. 12 presents the core prompts for updating the temporal strategy using the drift intensity and semantic revision signal.
Refer to caption Introspector: Generate Semantic Revision Signal You are [agent_name]. You can observe [observable_agent_name]. The total system reward for this time step is xx% lower than in the last time step indicating that the system state getting worse. Your individual reward has dropped from $x$ to $x$, which drop xx%.
@@
Step 2: According to the related rules you summarized, be specific about ‘‘what to do’’ and ‘‘what to avoid’’ and output between tags <self-checking_rules> and </self-checking_rules>.
@@
Figure 11: The core prompt of generating semantic revision signal.
@@
Based on the <reasons></reasons> you summarized (from Fig. 11), update your strategies carefully with NO MORE than xx%.
@@
[self-checking_rules] are a set of rules for verifying the output format and common-sense consistency of LLM outputs. These rules ensure that the model correctly applies physical reasoning. For example, in CPP task, “If ACCELERATION_GAP > 0, MY_VELOCITY will be INCREASED, MY_HEADWAY will be DECREASED.” These rules help correct occasional LLM mistakes, such as misjudging the travel direction of vehicles and thus misunderstanding how acceleration or deceleration affects inter-vehicle headway.
Appendix R AI Assistants in Writing
We utilized AI assistants (e.g., Gemini and Claude) exclusively for grammatical error correction, sentence polishing to enhance readability, and generating robot icons used in Figs. 1 and 2. We explicitly state that the core scientific concepts, experimental design, logical structure, reference, and substantive content of the paper were entirely human-authored without generative contributions from LLMs.